# Studying Workarounds in Software Forms: An Experimental Protocol

### v1

@@ Abstract
Workarounds enable users to achieve goals despite system limitations but expose design flaws, reduce productivity, risk compromising data quality, and cause inconsistencies. We propose an experimental method to investigate how users employ workarounds when the data they want to enter does not align with software form constraints. By conducting user studies based on this method, we can analyze how workarounds originate and impact system design and data integrity. Understanding workarounds is essential for software designers to identify unmet user needs.
@@
Workarounds Experimental Protocol User Study
@@ 1 Introduction
In Information Systems, software applications often misalign with user needs and goals [15], leading them to improvise solutions. These improvised workarounds are deliberate deviations from intended software usage that emerge when application forms and workflows fail to accommodate user needs [2]. Sometimes, developers recognize workarounds as useful system features and integrate them in their designs.
The presence of workarounds can have positive and negative consequences [3]. Workarounds can reveal missing features, overlooked design flaws, or inefficient workflows. For example, users might develop complex external processes to generate a report, unaware of an existing built-in functionality for the task. While workarounds can be beneficial by enabling users to overcome system limitations, they also pose risks. Deliberate deviations may lead to unintended consequences, such as data loss, system inconsistencies, or compromised data quality. For instance, a user entering irrelevant data to bypass a required field might unintentionally affect the accuracy of downstream reports, resulting in semantic inconsistencies where the system operates technically correctly but the data becomes unreliable [27]. Understanding the implications of workarounds is thus critical for improving software design and usability.
We propose an experimental method for investigating how users develop workarounds when their goals do not align with software constraints. This method enables descriptive analysis of workaround origins and their effects on system design and data integrity. Our approach provides insights into user behaviour during misalignments, advancing research in workaround detection and user-centered design. Analysis of these workarounds can inform automated detection mechanisms and recommendation systems for stakeholders in software design and usage.
@@ 2 Background and Related Work
3 Studying Workarounds
We aim to analyze the workarounds users do when interacting with forms in the context of misalignment between data and form software interfaces. We want to characterize their impact on user effort and data quality, with respect to frequency, and systematic behaviour.
By designing experiments using our proposed experimental method, researchers can investigate workarounds by examining their types, frequency, and relationships with each other and various types of misalignments.
Another important aspect to explore is the association between form widgets and the workarounds users attempt with them. This helps characterize different workaround strategies based on the available widgets users interact with when navigating software forms or workflows.
Additionally, researchers can analyze recurring patterns and systematic approaches users adopt when implementing workarounds. Another possible investigation is whether the frequency of misalignments affects the effort users invest in finding workarounds.
@@ 3.0.1 Experimental instruments
When investigating workarounds, researchers must clarify three key factors before conducting experiments: the creation of forms used in the study, the input data, and the assignment of users. Another critical instrument in workaround studies is a logger integrated into forms and workflows to capture user interactions, including widget usage, timing information, keystrokes, and mouse clicks. These logs are essential for analyzing interactions with form widgets and application navigation during workarounds.
For controlled or pilot experiments, researchers can use synthetic forms, data, and scenarios. They are responsible for recruiting and assigning users while designing various types of data misalignment, where input data does not fit the provided forms and workflows. Each form dataset should include both easily enterable records (where data format and content match expectations) and misaligned records.
@@
There are multiple form widget categorizations, such as “action”, “data”, and “static” [11] or “navigation”, “input”, “control” and “presentation” [12]. We propose a categorization based on a widget’s affordance for user workarounds, i.e., the degree of freedom it offers for data manipulation. Open: Widgets that allow users to freely modify input data, such as text boxes and text areas with minimal or no validation rules. Predetermined: Widgets containing preset data, requiring users to select from existing options, such as dropdown lists, combo boxes, radio buttons, checkboxes, and colour pickers. Controlled: Widgets that accept data within a constrained (but potentially large) range, such as sliders, number spinners, file uploads, and date pickers. When designing forms, researchers should balance the distribution of widgets from these categories to fit their controlled experiment.
Although the results of such experiments may not be fully generalizable, they provide valuable insights and build confidence in conducting further empirical studies, such as case studies or natural experiments. In a case study, researchers do not create the forms but advise the case study partner on software form and workflow design. There is no synthetic data, but researchers may provide guidance on user assignment. In a natural experiment, researchers do not influence form design or user assignment. Their role is limited to integrating the logger, which captures user interactions within the software forms and workflows.
These studies contribute to understanding workarounds and improving their detection, ultimately helping to identify missing requirements, validate software forms and workflows, and even automate pull request generation for missing features.
@@ 3.0.2 Experiment Process
In controlled experiments and case studies, each participant requires an individual session to enter their data into the system. Sessions consist of three parts: introduction, tasks, and exit survey. During the introduction, researchers present experimental tasks and provide documentation about the software forms or workflows, including screenshots of completed forms to familiarize participants with the interface and key features. Participants should remain unaware of the experiment’s focus on workarounds and receive instructions only to enter data through the forms and workflows. While loggers capture quantitative data, they may miss certain behaviours, necessitating the exit survey. After completing data entry tasks, participants must complete this survey about their experience. The survey can include questions about the participant’s occupation or field of study, self-assessment of difficulty entering misaligned data, self-reporting of workaround usage, and open-ended questions explaining employed workarounds or suggesting improvements to handle misalignments.
@@ 3.0.3 Metrics
In controlled experiments and case studies where misalignments and widget forms are known, researchers must label records that exposed a misalignment (i.e., for which a workaround is expected) according to the workaround type and quality. They also collect the widget involved and its widget category (open, predetermined, or controlled).
Workaround type reflects how users adapt to misalignments [28], while workaround quality assesses data integrity and semantic consistency between original and entered data [27]. Quality scores range from 1 to 4: A score of 1 is significant data loss, a score of 2 is partial data loss, a score of 3 is indirect data retention (entered elsewhere or reported differently while preserving intent), a score of 4 is prominent input of misaligned data, and no score when misaligned data completely skipped (no workaround).
@@
From record save times captured in the logger, researchers can deduce workaround implementation time. Since participants have unique data entry paces, researchers must calculate a distribution-sensitive baseline for normal record save times for each participant. After testing the distribution using Shapiro-Wilk: For a normal distribution, they can use the mean; for left/right-skewed distributions, the 30th/70th percentile, respectively; for heavy-tailed distributions, the trimmed mean (excluding the lowest/highest 10%); and for other cases, they can use the median.
@@
Another suggested metric is relative workaround effort which represents time to complete a workaround adjusted to baseline, computed as the ratio of workaround save time to participant’s baseline.

### v2

# How Users Employ Workarounds in Software Forms

@@ Abstract
Workarounds enable users to achieve goals despite system limitations but expose design flaws, reduce productivity, risk compromising data quality, and cause inconsistencies. This study investigates how users employ workarounds when the data they want to enter does not align with software form constraints. Through a descriptive user study, we analyzed how workarounds originate and impact system design and data integrity. Understanding workarounds is essential for software designers to identify unmet user needs.
@@
Workarounds User study Software application
@@ 1 Introduction
In Information Systems, software applications often misalign with user needs and goals [15], leading them to improvise solutions. These improvised workarounds are deliberate deviations from intended software usage that emerge when application forms and workflows fail to accommodate user needs [2]. Sometimes, developers recognize workarounds as useful system features and integrate them in their designs. For instance, developers designing a class diagram in Ecore within the Eclipse Modeling Framework often specify arbitrary values such as -1 or 1000 in the upper bound of an association to denote an unlimited number of targets, rather than using the UML-prescribed “*”. This workaround has become institutionalized as a documented feature in Eclipse because the cardinality text box only supports integer values.
The presence of workarounds can have positive and negative consequences [3]. Workarounds can reveal missing features, overlooked design flaws, or inefficient workflows. For example, users might develop complex external processes to generate a report, unaware of an existing built-in functionality for the task. While workarounds can be beneficial by enabling users to overcome system limitations, they also pose risks. Deliberate deviations may lead to unintended consequences, such as data loss, system inconsistencies, or compromised data quality. For instance, a user entering irrelevant data to bypass a required field might unintentionally affect the accuracy of downstream reports, resulting in semantic inconsistencies where the system operates technically correctly but the data becomes unreliable [28].
Understanding the implications of workarounds is thus critical for improving software design and usability. This is especially the case for low-code applications, whose interfaces are typically forms and pages [18]. There is a lack of research on the workarounds employed by software engineers [14], and therefore by extension, citizen-developers [7] in low-code applications. We envision that insights into workarounds can be used to guide developers and product owners in identifying opportunities for new features, refine validation mechanisms, and enhance software maintenance. For example, recognizing frequent use of comment boxes for unsupported data could lead to redesigning the form to better accommodate such information.
Here, we investigate how users handle misaligned cases where user goals or data requirements do not align with the constraints of software forms or workflows. Through a descriptive analysis, we examine the origins of workarounds [20] and their broader implications for system design and data integrity. The study provides insights into user behaviour when encountering misalignments. We contribute a publicly available dataset of user workarounds and present key findings to advance research on workaround detection and user-centered software design. Analyzing workarounds can inform future automatic detection mechanisms and recommendation systems for stakeholders in software design and usage.
We discuss related work in Section 2, and outline the user study in Section 3. We present our findings in Section 4, and discuss them in Section 5. We mention threats to validity in Section 6 and conclude in Section 7.
@@ 2 Background and Related Work
3 Experiment Design
We aim to analyze the workarounds users do when interacting with forms in the context of misalignment between data and form software interfaces. We want to characterize their impact on user effort and data quality, with respect to frequency, and systematic behaviour. We pose the following research questions:
  • RQ1.
    What are the characteristics of workarounds users employ in software application forms?
    • RQ1.1.
      Which types of workarounds do users employ, if any?
    • RQ1.2.
      How frequently do users employ workarounds ?
    • RQ1.3.
      Which types of workarounds are most commonly associated with data misalignment types?
    We thus categorize workarounds, measure their frequency, and assess the quality of data input when users utilize workarounds.
  • RQ2.
    Which form widgets are most commonly associated with workarounds?
    We thus analyze patterns in the widget involved during workarounds.
  • RQ3.
    How do users handle recurring data misalignments?
    • RQ3.1.
      How does the frequency of data misalignments affect user effort in workarounds?
    • RQ3.2.
      To what extent do users develop systematic approaches when implementing workarounds?
    We therefore examine how repeated exposure to specific misalignments influences the time users spend on workarounds and investigate whether these workarounds are applied systematically or tailored to individual cases.
3.0.1 Setup
We conducted an experiment [27] to study user workarounds in software applications by analyzing interactions with forms for entering misaligned data. Participants used a Windows 11 PC (AMD Ryzen 7 5700U, 1.8GHz, 16GB RAM) with a 24-inch monitor and Python 3.11.7, and sessions were conducted via Microsoft Teams remote control. Participants were tasked with entering prepared datasets into instrumented software applications we developed. We extracted entered data, timing information, keystrokes and mouse clicks from the generated logs to analyze the interactions with form widgets and application navigation. We included cases where a participant entered and then corrected text.
@@ 3.0.2 Experimental instruments
We designed three simple applications with Python and CustomTkinter [^1] https://github.com/TomSchimansky/CustomTkinter 5.2.2: a grading app “TA” for teaching assistants tasked with entering student grades for an exam; a human resources app “HR” where HR professionals can complete an employee onboarding form; and an event registration app “REG” for clerks to register participants to a conference. All three consist of a main form that users can complete with records from the provided datasets. They also allow navigating to previously entered records and modifying them, submitting feedback in free-text globally for all the entered records, and reviewing all data before final submission.
@@
We created three versions of TA: Open, Predetermined, and Controlled, based on the different types of widgets used. Lo et al. classified widgets as “action”, “data”, and “static” [11]; Meneses et al. classified them as “navigation”, “input”, “control”, and “presentation” [12]. Our classification rather focuses on the affordance of widgets for user workarounds, i.e., the degree of freedom that a widget offers to a user to manipulate. The Open version consists solely of widgets that allow users to freely modify the input data they capture, such as text boxes and text areas with minimal or no validation rules. The Predetermined version consists solely of widgets that contain preset data, requiring users to select from existing option, such as dropdown lists, combo boxes, radio buttons, checkboxes, and colour pickers. The Controlled version consists solely of widgets that accept data within a constrained (but potentially large) range of values such as sliders, number spinners, file uploads, and date pickers. The HR and REG apps offer a mix of all these three widget categories.
For each app, we synthesized a dataset consisting of 26 records for participants to enter. Each dataset contains 11 records that can be easily entered in the application, i.e., the data format and content align exactly with what the main form expects. The remaining 15 records contain predefined data misalsignments. We define six misalignment types that share the same intent for each application. Since these types may overlap, we do not consider this separation as a partition but as a logical grouping of misalignments. Explicit: This misalignment is clearly visible before entering the record, e.g., in TA, the total is 105 (indicating that there was a bonus question), whereas normal grades are capped at 100. Implicit: This misalignment is not immediately visible, e.g., in TA, the record is marked to have completed the bonus, and the total grade is 95. Excess: The record contains more information than the form can capture, e.g., the widget for the meal preference in REG cannot accommodate the detailed dietary requirements provided in the record. Nonstandard: The data is formatted in an unconventional way, e.g., a website rating widget in REG requires scores on a scale from 1 to 5, but the dataset contains the rating as a percentage. Outlier: The data provided is nonsensical, e.g., in HR a record lists -1 hour for a widget capturing allowable remote work hours. Ambiguous: The data can be interpreted in multiple ways, e.g., it is not clear what multiple phone numbers within the same record mean and how to enter them in a single widget of HR.
Each dataset randomly includes 3 misalignment types with 5 instances of each type. The 26 records of each dataset are randomly ordered for each participant.
@@ 3.0.3 Experiment Process
Each participant had an individual session where they attempted to enter a provided dataset to the app. Each session had three parts: introduction, tasks, and exit survey. In the introduction, we presented the tasks and provided documentation about the apps. The document included screenshots of completed forms to familiarize participants with the user interface and key features. This part lasted 10 minutes.
We then asked participants to complete three data entry tasks, one with each app described above. The version for the TA app was assigned randomly but balanced across all participants. Participants were unaware of the goal of the experiment and were instructed only to interact with three apps to enter data, with no mention of workarounds. Although we did not enforce time limits for the experiment, participants took on average 1.5 hours to complete the experiment. After completing the data entry tasks, participants completed a survey consisting of 12 questions structured as follows: Background: 1 open-ended question about the participant’s primary occupation or field of study. Ease of data entry: 5 five-point Lickert-scale questions asking them to self-assess the ease and difficulty of entering the records. Workarounds: 5 five-point Lickert-scale questions on whether they resorted to workarounds for the misaligned records and on whether some workarounds were recurring. Explanation and feedback: 3 open-ended questions asking them to explain why some records were difficult to enter, to describe the workarounds they employed, and suggest how the forms could be adapted to handle the misalignments.
The experiment protocol, datasets, apps, exit survey, and statistical results are available in the replication package online.222 https://zenodo.org/records/14292103
3.0.4 Participants
We used convenience sampling, sending invitations to research groups and communities within our network. No prior knowledge of any specific software was required. We hired 17 volunteer for the study: two in a pilot round to calibrate the experiment and refine the analysis, and 15 participants completed the experiment. Among them, we count 11 graduate students (including the two for the pilot), one bachelor student, and 5 professionals. 14 had a background in computer science, while 3 held degrees in other engineering disciplines.
Figure 1: Distribution of the data misalignments across the 15 participants
Fig. 1 shows the distribution of the data misalignment types that each participant faced across the 3 tasks. Each participant encountered at least 4 misalignment types for a total 45 misaligned data. We analyzed how participants employed workarounds for 120 Implicit, Excess, Ambiguous Nonstandard, 100 Outlier, and 95 Explicit misalignments, for a total of 675 misaligned records.
@@ 3.0.5 Metrics
One author performed the initial labelling of the collected data, and another reviewed the labels. Conflicts were resolved collaboratively in a single session. We labelled each record that exposed a misalignment (i.e., for which a workaround is expected) from each participant according to the workaround type and quality. We also collected the widget involved when a workaround occurs and the widget category (open, predetermined, or controlled).
Workaround type reflects the ways users adapt to misalignments [29]. Data Adjustment (DA): modifying the original data to align with application constraints; Functional Adjustment (FA): using the form’s widgets or functionalities in unintended ways; Fallback (FB): resorting to an alternative, such as a general comment box, to address the misalignment; Data Loss (DL): truncating or omitting parts of the data, without any strategic approach; No Workaround (NW): abandoning any attempt to resolve the misalignment.
Workaround quality assesses the integrity of the entered data and how much it differs from the intent of the provided misaligned data. It estimates to what extent the original data and the entered data are semantically consistent [28] and whether a human or machine can make similar inferences from them. We give a score from 1 to 4 according to the following schema. A quality score of 1 indicates significant loss of the provided misaligned data compared to what the participant entered. A score of 2 reflects partial loss of the misaligned data. A score of 3 means the participant retained the misaligned data indirectly, such as entering it elsewhere in the app or reporting it in a way that preserves the original intent. A score of 4 is assigned when the participant manages to prominently input the misaligned data. If the misaligned data is completely skipped, it is considered as NW, with no quality score assigned.
@@
From the save time of each record, we deduced the time each participant took to implement a workaround for each misaligned record. Recall that, for each app, participants are provided with both normal and misaligned records. Since each participant has a unique pace when entering data, regardless of data alignment, we calculate a distribution-sensitive central tendency measure as the baseline for normal record save times for each participant in each app. To improve accuracy, we test the distribution of save times of each participant for each app using the Shapiro-Wilk test. If the distribution is normal, the baseline is the mean. For left-skewed or right-skewed distributions, the baseline is the 30th or 70th percentile respectively. For heavy-tailed distributions (based on kurtosis), we compute the trimmed mean, excluding the lowest and highest 10% of data points. In all other cases, the baseline is the median.
@@
Relative workaround effort represents the time a participant takes to complete a workaround adjusted to the baseline. It is computed for each misaligned record as the ratio of the save time when completing the workaround and the participant’s baseline for that app.